### v2

# Safe-CLIP: Removing NSFW Concepts from Vision-and-Language Models

Abstract
Large-scale vision-and-language models, such as CLIP, are typically trained on web-scale data, which can introduce inappropriate content and lead to the development of unsafe and biased behavior. This, in turn, hampers their applicability in sensitive and trustworthy contexts and could raise significant concerns in their adoption. Our research introduces a novel approach to enhancing the safety of vision-and-language models by diminishing their sensitivity to NSFW (not safe for work) inputs. In particular, our methodology seeks to sever “toxic” linguistic and visual concepts, unlearning the linkage between unsafe linguistic or visual items and unsafe regions of the embedding space. We show how this can be done by fine-tuning a CLIP model on synthetic data obtained from a large language model trained to convert between safe and unsafe sentences, and a text-to-image generator. We conduct extensive experiments on the resulting embedding space for cross-modal retrieval, text-to-image, and image-to-text generation, where we show that our model can be remarkably employed with pre-trained generative models. Our source code and trained models are available at: https://github.com/aimagelab/safe-clip. $*{}^{*}$qual contribution.
Keywords:
Trustworthy AI Vision-and-Language NSFW Concepts
Warning: This paper includes explicit sexual content, racially insensitive language, and other material that may be disturbing or offensive to certain readers.
1 Introduction
Large-scale models have recently proven to be effective on a variety of tasks, ranging from image classification and understanding to cross-modal retrieval and generation [37, 40, 30]. Scaling models, however, has also required to increase the quantity and variability of training data, paving the way to scraping billions of items from the web without manual supervision [45, 44]. Despite the adoption of filters and automatic checks, this paradigm comes at the cost of introducing inappropriate content in the training set [42, 15], which ultimately results in the injection of unsafe, biased or toxic behaviors [4].
This is also the case of vision-and-language models based on embedding spaces, where toxic content can embed itself in the latent space. For instance when a NSFW (not safe for work) textual prompt is used for a cross-modal task, its embedding can reach unsafe points in the latent space, leading to the generation of undesired images, or to the retrieval of inappropriate content. Similarly in image-to-text generation, when an inappropriate image is used as a prompt, the descriptive text could be toxic or offensive. A qualitative example of this is reported in Fig. 1, considering the case of a CLIP backbone [37].
Figure 1: Removing NSFW concepts from CLIP models. Our Safe-CLIP fine-tunes CLIP to make it safer in cross-modal retrieval, image-to-text and text-to-image generation.
Driven by these considerations, we tackle the task of enhancing the safety of pre-trained vision-and-language models. In particular, we devise an approach for making a CLIP-style embedding space safer, so that it becomes invariant with respect to inappropriate inputs. While some previous attempts have focused on mitigating inappropriate content in diffusion models [42], our approach mitigates inappropriate concepts from CLIP-like embedding spaces. As such, it has a more general impact and applicability, as CLIP-like models are employed for many different applications, ranging from cross-modal retrieval [37] to text-to-image and image-to-text generation [40, 30], and are employed as feature extractors for different tasks [46, 51, 33].
Our approach is based on fine-tuning the embedding space so as to avoid the representation of inappropriate content, without changing its normal expressive power. We do this with a combination of losses designed to redirect inappropriate content to safe embedding regions, while preserving the structure of the embedding space as intact as possible. To support our training procedure, we generate quadruplets of safe and unsafe vision-language items. This data generation strategy is empowered by a toxic language model which can translate safe textual prompts into unsafe ones, while keeping context alignment and semantic meaning unchanged. When applied to collections of visually-grounded descriptions in conjunction with an NSFW-capable diffusion model, our conditioned NSFW generator is employed for building a dataset which can properly support the fine-tuning of a CLIP embedding space.
The resulting safe version of CLIP can be applied to cross-modal retrieval, text-to-image and image-to-text generation (Fig. 1). For example, if we ask the fine-tuned version of CLIP to retrieve an image corresponding to a textual prompt with NSFW content, it will fetch an image with similar semantics but appropriate content. Also, a Stable Diffusion model [40] conditioned on our fine-tuned CLIP will generate an image with appropriate content, free of violence, nudity, or other toxic aspects, keeping the safe semantics of the input prompt. Similarly, a multimodal LLM like LLaVA [30] conditioned on our Safe-CLIP will generate a textual description without inappropriate content.
Experimentally, we evaluate the capabilities of our strategy for making CLIP safer in both retrieval and generation contexts, by running experiments both on prompts and images synthetically generated and employing existing unsafe prompts [42] and real images. Experimental results show that our approach can significantly improve the safety during text-to-image and image-to-text retrieval and during visual and textual generation.
To sum up, our main contributions are as follows:
  • We introduce a novel fine-tuning methodology which can turn a pre-trained CLIP-like embedding space into a safer one. Once fine-tuned with our methodology, the CLIP space ignores NSFW content and can be applied to downstream tasks like retrieval and visual or textual generation.
  • Our approach is based on creating a toxic LLM which can generate unsafe prompts given safe and visually-grounded ones. This is obtained by fine-tuning Llama 2 on manually curated pairs, and then aligning it with Direct Preference Optimization (DPO).
  • Leveraging an automatically generated dataset of safe and unsafe images and texts, we fine-tune CLIP with a novel combination of losses which redirect unsafe content while preserving the embedding space structure.
  • We experimentally evaluate the appropriateness of our approach by conducting experiments in both retrieval, and textual and visual generation. Our method showcases a significantly reduced generation of NSFW content.
2 Related Work
Removing concepts from vision-and-language models. Removing content from AI models has been recently gaining increasing attention, with techniques spanning from complete model retraining or fine-tuning to machine unlearning [6, 17, 18] and differential privacy [19]. Some of these attempts have been considering text-to-image models and have aimed at deleting styles, concepts, or objects [25, 54]. Recently, Schramowski et al. [42] introduced a technique to steer the generation away from NSFW areas, defined by a finite and fixed set of concepts. NSFW concepts are encoded with the input prompt at inference time and the NSFW embedding is used as negative guidance. Later, Gandikota et al. [15] proposed a fine-tuning method that can erase a visual concept given only its name and using negative guidance as a teacher.
In contrast to these previous works, we focus on removing NSFW from a contrastive CLIP-like model, which can be applied for cross-modal retrieval, and for visual and textual generation. While to the best of our knowledge we are the first to tackle this scenario, Trager et al. [49] have demonstrated the presence of compositional patterns within the embedding space of CLIP, which suggests the existence of a distinctive path from safe to NSFW zones.
Detecting NSFW content. A related research field is that of the automatic detection of NSFW content. Several approaches have been proposed to detect NSFW language [7, 20, 32], primarily on social media data sources. DistilBERT [41] emerges as a promising solution for this purpose, particularly when fine-tuned for adult content detection. We utilize it as an NSFW language detector, in conjunction with GPT-3.5 [36], which we query directly to classify our prompts. While the identification of unsafe language poses a challenging task, the same can be said for vision, where different approaches have been proposed to detect inappropriate content [14, 34, 3]. Still, the detection of inappropriate content remains an intricate challenge, as visual cues, lack of context, and restricted data sources often introduce added layers of complexity. In our analysis, we utilize Q16 [43] and NudeNet [2] as automatic detectors of NSFW images. In particular, NudeNet is specialized in identifying unsafe content related to nudity, while Q16 serves as a broader spectrum NSFW classifier.
Finetuning LLMs with little data. Large Language Models (LLMs) have achieved high performance in various tasks due to their zero-shot capabilities [38, 47, 48], which stems from model scaling and the utilization of large training datasets. In addition to fine-tuning these models for specific tasks [8, 56, 31], there has been recently an interest in building parameter efficient fine-tuning strategies. In most solutions, only part of the weights are trained [55, 16] or a reduced number of weights are added to the LLM [11, 21]. As shown in [52], datasets employed for supervised fine-tuning [1] play a central role in changing the LLM behavior [57], even in low-data regimes. In this work, we fine-tune Llama 2 [48] to produce unsafe prompts starting from pre-existing safe counterparts.
3 Proposed Approach
CLIP-like models [37] are trained on web-crawled data which can contain inappropriate content [4]. Making these models safer, therefore, requires either retraining from scratch using large-scale cleaned data or fine-tuning them with a form of supervision that aims to mitigate inappropriate knowledge. The first option would necessitate data cleaning at large scale, which is currently not effective in practice (see also Sec. 4.3 for a comparison), so we instead employ the second strategy. Specifically, we focus on making both the textual encoder and the visual encoder of CLIP safer.
Ideally, we want a safe version of the CLIP text encoder to ignore inappropriate content from input sentences and understand most of its clean content. Symmetrically, we want the safe version of the CLIP visual encoder to ignore inappropriate content from input images. Furthermore, we also want to maintain as much as possible the original structure of the embedding space near safe textual or visual regions, so that the safe encoders can be straightforwardly connected to downstream models built on top of them without further adaptation. Formally, given an unsafe sentence $t_{i}^{\star}$ and a “cleaning” function $c_{t}(\cdot)$ which removes all inappropriate content from it, we want our safe textual encoder $\mathcal{T}$ to satisfy the following condition with respect to the original, pre-trained, CLIP text encoder $\mathcal{T}_{0}$:
$\mathcal{T}(t_{i}^{\star})\approx\mathcal{T}(c_{t}(t_{i}^{\star}))\approx\mathcal{T}_{0}(c_{t}(t_{i}^{\star})),$ (1)
where with the $\approx$ sign we indicate high similarity in the embedding space. As it can be noticed, the first condition stated in Eq. 1 ensures that inappropriate content is ignored, while the second provides that the safe CLIP textual encoder can properly encode the cleaned part of the input sentence. On the other hand, this also ensures that $\mathcal{T}$ can be seamlessly connected to downstream models that were trained on the basis of $\mathcal{T}_{0}$ (for instance, Stable Diffusion v1.4 [40] in the case of a CLIP ViT-L/14). The same requirement is applied to the visual encoder: given an unsafe image $v_{i}^{\star}$ and a visual “cleaning” function $c_{v}$, we require:
$\mathcal{V}(v_{i}^{\star})\approx\mathcal{V}(c_{v}(v_{i}^{\star}))\approx\mathcal{V}_{0}(c_{v}(v_{i}^{\star})),$ (2)
where $\mathcal{V}$ is the safe visual encoder and $\mathcal{V}_{0}$ is the original CLIP visual encoder.
3.1 Building the ViSU dataset
Overview. In order to modify the CLIP model to avoid the representation of inappropriate content, our methodology requires a dataset comprising quadruplets of safe and unsafe (i.e. NSFW) images and sentences, denoted as:
$\mathcal{D}=\{(v_{i},t_{i},v_{i}^{\star},t_{i}^{\star}),i=1,...,N\},$ (3)
where $v_{i}$ indicates a safe image, $t_{i}$ its corresponding sentence, and the unsafe image $v_{i}^{\star}$ and unsafe sentence $t_{i}^{\star}$ are “paired” to convey a similar semantic meaning of their safe counterparts. For instance, $t_{i}$ can be considered as the sanitized version of $t_{i}^{\star}$, expressing a similar meaning without inappropriate concepts, and the same holds for their visual counterparts. As such a dataset is not available, we build $\mathcal{D}$ with an automatic annotation procedure where 1⃝ unsafe sentences $t_{i}^{\star}$ are automatically generated starting from cleaned sentences $t_{i}$, and 2⃝ unsafe images $v_{i}^{\star}$ are generated starting from unsafe sentences $t_{i}^{\star}$.
Training a conditioned NSFW textual generator. To achieve the first goal, we fine-tune a large language model (Llama 2-Chat [47]) to generate unsafe sentences starting from safe ones. In particular, we employ a set of 100 manually-curated safe-unsafe pairs, building these as a mixture of manually written pairs and sentences generated automatically with Vicuna [8]. To ensure that the dataset provides proper supervision, we follow the definition of NSFW content of [42] as that of content belonging to the following twenty categories: hate, harassment, violence, suffering, humiliation, harm, suicide, sexual, nudity, bodily fluids, blood, obscene gestures, illegal activity, drug use, theft, vandalism, weapons, child abuse, brutality, cruelty, and balance the samples of the training dataset across these categories to encourage the LLM to generate unsafe content with good variety.
Figure 2: Overview of our Safe-CLIP approach.
[…]
Aligning the textual NSFW generator. With the aim of increasing the quality of generated unsafe sentences, and also their semantic relatedness to the prompt, we adopt a fine-tuning stage by devising a variant of Direct Preference Optimization (DPO) [39]. DPO was originally proposed as an alternative to RLHF [9] with better stability and which does not need the explicit training of a reward model. Like RLHF, however, DPO employs large-scale human preference annotations, which in our case are not available. As a replacement, we build an automatic ranking procedure which can replace the human preference annotation while still increasing the alignment of our NSFW generator.
In particular, given a safe text $t_{i}$, we obtain two different unsafe completions $(t_{i,0}^{\star},t_{i,1}^{\star})$ from the SFT model by sampling from its output probability distribution. We then rank the quality of the obtained completions by considering their NSFW degree and their semantic similarity with $t_{i}$. For the former criterion, we obtain a binary NSFW rating $\text{nsfw}(t_{i}^{\star})\in\{0,1\}$ by prompting GPT-3.5 with a completion. For the latter, we instead employ the CLIP similarity between $t_{i}$ and each of the completions, as predicted by the pre-trained text encoder using cosine similarity, thus lying in the range $\left[-1,1\right]$. The final quality degree of an unsafe completion $t_{i}^{\star}$, given its safe prompt $t_{i}$, is then obtained as
$\text{rank}(t_{i}^{\star},t_{i})=\text{CLIP-Sim}(t_{i}^{\star},t_{i})+\text{NSFWRate}(t_{i}^{\star}),$ (4)
[…]
Overall, our SFT and preference optimization pipeline turns Llama 2-Chat into a powerful generator of textual NSFW content, which can also perfectly maintain semantic relatedness with respect to a safe input sentence. What is more, our NSFW generator can still support diverse prompts, different from those seen at training time. Interestingly, for instance, it can be asked to add NSFW content belonging to a specific category (e.g. violence or nudity).
Generating the full ViSU dataset. Having a conditioned NSFW generator, we can generate NSFW texts starting from safe, visually relevant, sentences. Starting from NSFW sentences, we then generate corresponding NSFW images $v_{i}^{\star}$ using a diffusion-based model which has been trained on NSFW content [^2] Specifically, we use the stablediffusionapi/newrealityxl-global-nsfw model available on HuggingFace, which has a high probability of generating NSFW images (i.e. around 80% of images have been detected as NSFW in our experiments).. The overall dataset, which we term ViSU (Visual Safe-Unsafe), contains 165k quadruplets of safe and unsafe sentences and images generated starting from COCO Captions [27]. We follow the Karpathy’s splits [22] to organize the dataset into training, validation and test splits. To ensure that the resulting dataset is balanced across the twenty NSFW categories, we prompt the NSFW generator by asking to inject a specific, randomly chosen, category into the safe sentence. Sample quadruplets from our dataset are reported in the supplementary.
3.2 Making CLIP Safe
Having built a dataset with quadruplets $(v_{i},t_{i},v_{i}^{\star},t_{i}^{\star})$ of safe and unsafe images and sentences, we make the CLIP model safe with a procedure that ensures that the conditions expressed in Eq. 1 and 2 are met. To this aim, we adopt a multi-modal training scheme with four loss functions. Specifically, we define two inappropriate content redirection losses that aim at teaching the model to ignore unsafe content in an input text or input image, and two structure preservation losses that aim at maintaining the original structure of the embedding space in safe regions. During the rest of this section, $\mathcal{T}$ and $\mathcal{V}$ will indicate the textual and visual encoders being fine-tuned, $\mathcal{T}_{0}$ and $\mathcal{V}_{0}$ frozen “deep copies” of the textual and visual encoders obtained before the fine-tuning starts.
Inappropriate content redirection. To teach the model to ignore inappropriate content, we propose to impose cross-modal similarities between unsafe sentences $t_{i}^{\star}$ and corresponding images $v_{i}$ in the dataset, and between unsafe images $v_{i}^{\star}$ and corresponding texts $t_{i}$. Noticeably, this is not granted in $\mathcal{T}_{0}$ and $\mathcal{V}_{0}$, which instead will have good metric learning properties between $t_{i}$ and $v_{i}$. To encourage this effect even further, we also require that the embedding of the unsafe sentence $t_{i}^{\star}$ matches that of the corresponding safe sentence $t_{i}$ according to the frozen textual encoder, and the embedding of unsafe images $v_{i}^{\star}$ matches that of the corresponding safe images $v_{i}$ according to the frozen visual encoder, through a cosine similarity term which only considers positive pairs and ignore distances with respect to negative samples.
Formally, given a batch of $N$ images $\mathbf{V}=\left[v_{1},v_{2},...,v_{N}\right]$, their corresponding safe captions $\mathbf{T}=\left[t_{1},t_{2},...,t_{N}\right]$, unsafe texts $\mathbf{T}^{\star}=\left[t_{1}^{\star},t_{2}^{\star},...,t_{N}^{\star}\right]$ and unsafe images $\mathbf{V}^{\star}=\left[v_{1}^{\star},v_{2}^{\star},...,v_{N}^{\star}\right]$, we define two $N\times N$ matrices containing pairwise cosine similarities between $\mathbf{T}^{\star}$ and $\mathbf{V}$ and between $\mathbf{V}^{\star}$ and $\mathbf{T}$. We then adopt a symmetric InfoNCE loss [35] which aims at maximizing the cosine similarity between the $N$ matching pairs of cross-modal safe and unsafe embeddings, and minimize those of the $N^{2}-N$ non-matching pairs while having, in turn, one of the encoders frozen and the other being fine-tuned:
[…]
where $\tau$ is a temperature parameter. The second loss term, which brings each unsafe sentence close to its corresponding safe one, and each unsafe image close to its corresponding safe one, is instead expressed as the negative cosine similarity between each unsafe embedding and the original safe embeddings, as follows:
$L_{\text{redir,2}}=-\frac{1}{N}\left(\sum_{i=1}^{N}\cos(\mathcal{T}(t_{i}^{\star}),\mathcal{T}_{0}(t_{i}))+\sum_{i=1}^{N}\cos(\mathcal{V}(v_{i}^{\star}),\mathcal{V}_{0}(v_{i}))\right).$ (6)
Embedding structure preservation. The two aforementioned losses bring unsafe embeddings towards the positions of their corresponding safe embeddings in the original frozen spaces, either in a single-modal or multi-modal manner. However, alone, they would inevitably cause the fine-tuned encoders $\mathcal{T}$ and $\mathcal{V}$ to lose their performance on safe inputs, as well as their matching properties. Therefore, we also adopt two losses to ensure that the original structure of the embedding spaces is preserved where safe textual and visual regions lie.
In particular, we define a matching loss between the safe embeddings produced by the online networks $\mathcal{T}$, $\mathcal{V}$ and those of the original, pre-trained, networks $\mathcal{T}_{0}$, $\mathcal{V}_{0}$. Again, this is defined through the negative cosine similarity between matching pairs, as
$L_{\text{pres,1}}=-\frac{1}{N}\left(\sum_{i=1}^{N}\cos(\mathcal{T}(t_{i}),\mathcal{T}_{0}(t_{i}))+\sum_{i=1}^{N}\cos(\mathcal{V}(v_{i}),\mathcal{V}_{0}(v_{i}))\right).$ (7)
Finally, as an additional regularization term we also keep a contrastive loss between safe visual embeddings and safe textual embeddings, comparing on-line and frozen encoders. This, in practice, closely resembles the original loss on which the embedding space was trained, i.e.
[…]
Eventually, the overall loss function on which the network is trained is a weighted sum of the four loss functions mentioned above.
4 Experiments
4.1 Experimental Setting
Datasets. Our experiments are mainly conducted on the collected ViSU dataset that, as previously mentioned, contains 165k quadruplets of safe and unsafe content for training. For both validation and testing we use 5k samples following the Karpathy’s COCO splits [22], randomly sampling only one safe caption among the five available for each image in the original COCO dataset. When applying our Safe-CLIP to text-to-image generative architectures, we also perform experiments on the I2P dataset [42] that is composed of 4,703 textual prompts extracted from Lexica [^3] https://lexica.art/, a collection of user-generated prompts for conditioning text-to-image diffusion models. Each prompt is associated to one of seven different categories of inappropriate content, among hate, harassment, violence, self-harm, sexual content, shocking images, illegal activity.
LLM fine-tuning details. During SFT, we fine-tune the 7B version of Llama 2-Chat using low-rank adaptation [21] with $r=64$ as low-rank factor. We employ a batch size equal to 4 and a learning rate of $2\times 10^{-4}$. To perform DPO, we follow the variant presented in [50], employing low-rank adaptation also in this case. The complete DPO fine-tuning settings are reported in the supplementary.
Safe-CLIP implementation and training details. Our architecture is based on the standard CLIP model [37] composed of a visual and a textual encoder. Specifically, we employ the ViT-L/14 variant to comply with the textual encoder used in the Stable Diffusion v1.4 model [40] and the visual encoder employed in LLaVA [30]. Experiments with different CLIP-based backbones are reported in the supplementary. During training, both the visual and textual encoder are fine-tuned using low-rank decompositions [21], where the low-rank factor $r$ is set to $16$ in all the experiments. We employ Adam as optimizer [24] using a learning rate equal to $1\times 10^{-3}$ and a batch size of 128.
4.2 Evaluating the ViSU Dataset
Table 1: Comparison between the textual portion of ViSU and the I2P benchmark [42], in terms of NSFW degree and toxicity.
% NSFW
Dataset DistilBERT GPT-3.5 Toxicity
I2P [42] 52.8 13.9 14.9
w/o SFT (i.e. Llama 2-Chat) 37.8 9.3 7.7
w/o DPO fine-tuning 75.9 75.0 30.6
ViSU (Ours) 80.9 79.1 31.3
We first assess the quality of our ViSU dataset, evaluating the inappropriateness degree of the generated unsafe sentences. Specifically, we employ a DistilBERT model fine-tuned for adult content detection and directly ask GPT-3.5 to evaluate whether the generated unsafe sentences should be classified as NSFW content. As reported in Table 1, ViSU showcases a very good textual quality, as it has a higher degree of NSFW sentences compared to existing alternatives like I2P [42] (i.e. 79.1% of NSFW sentences vs. 13.9% in the I2P dataset according to GPT-3.5). Moreover, following [42], we also report the toxicity score of the unsafe sentences computed using the Perspective API [^4] https://github.com/conversationai/perspectiveapi. Also in terms of toxicity, ViSU presents a higher degree of NSFW content. When instead comparing the effectiveness of each of the Llama 2 fine-tuning stages, it is worth noting that both the SFT procedure and DPO fine-tuning consistently increase the quality of generated sentences, going from 9.3% of NSFW content when using the original Llama 2 model to 79.1% after both fine-tuning stages.
4.3 Evaluating the Safe-CLIP Embedding Space
Results on ViSU test set. To evaluate the retrieval performance of Safe-CLIP, we firstly consider image-to-text and text-to-image retrieval in a safe-only setting, where we do not have any inappropriate content in both visual and textual data. This is important to assess whether the properties of the original CLIP embedding space are preserved when employing our fine-tuning strategy. In this case, query elements are represented by the safe images of the test set (which, with a slight abuse of notation, we refer to as $\mathbf{V}$) for the image-to-text setting and the safe textual items (referred to as $\mathbf{T}$) for the text-to-image one. Moreover, we consider text-to-image and image-to-text retrieval when using unsafe texts as queries (referred to as $\mathbf{T}^{\star}$) and both safe and unsafe images as retrievable items and when using unsafe images as queries (i.e. $\mathbf{V}^{\star}$) and both safe and unsafe texts as retrievable items.
Table 2: Retrieval results on the ViSU test set. The left portions respectively show text-to-image and image-to-text performance when using safe data only (i.e. $\mathbf{V}$ and $\mathbf{T}$). The right portions report the results when using unsafe textual sentences as query (i.e. $\mathbf{T}^{\star}$) and the merging of safe (i.e. $\mathbf{V}$) and unsafe images (i.e. $\mathbf{V}^{\star}$) as retrievable items, or when using unsafe visual queries (i.e. $\mathbf{V}^{\star}$) and the merging of safe (i.e. $\mathbf{T}$) and unsafe sentences (i.e. $\mathbf{T}^{\star}$) as retrievable items.
Text-to-Image Image-to-Text Text-to-Image Image-to-Text
($\mathbf{T}$-to-$\mathbf{V}$) ($\mathbf{V}$-to-$\mathbf{T}$) ($\mathbf{T}^{\star}$-to-$\mathbf{V}\cup\mathbf{V}^{\star}$) ($\mathbf{V}^{\star}$-to-$\mathbf{T}\cup\mathbf{T}^{\star}$)
Model R@1 R@10 R@20 R@1 R@10 R@20 R@1 R@10 R@20 R@1 R@10 R@20
CLIP (ViT-L) [37] 36.8 71.6 81.5 39.8 74.2 83.5 2.0 24.8 33.2 4.5 32.9 40.6
DataComp-1B (ViT-L) [13] 46.7 79.7 87.4 47.0 81.3 88.9 1.6 28.1 35.6 5.5 37.5 44.9
w/o inap. content redirection 49.9 83.7 90.3 48.1 83.6 90.5 1.6 30.4 40.1 6.1 35.2 42.6
w/o negative cosine similarities 41.9 78.5 87.3 41.5 77.8 86.9 8.2 46.0 56.6 13.7 60.4 68.2
Safe-CLIP 45.9 81.8 89.7 45.3 82.3 89.7 8.0 46.9 58.0 19.1 62.9 71.1
Retrieval results on the ViSU test set are reported in Table 2, comparing the proposed Safe-CLIP model with the original CLIP architecture, a CLIP model trained on the DataComp dataset [13], which has undergone NSFW content cleaning, and two different baselines. Specifically, we consider a variant of our approach in which we remove the two negative cosine similarity losses (i.e. Eq. 6 and 7), and a model trained with safe data only (i.e. removing the loss functions for inappropriate content redirection reported in Eq. 5 and 6). Results are reported in terms of Recall@$k$ (R@$k$) with $k=1,10,20$, that measures the percentage of times the visual or textual item associated to the query is retrieved among the top-$k$ elements. When using unsafe sentences as queries, for each element we consider the safe image associated with the given text as the corresponding visual element. Symmetrically, when using unsafe images as queries, for each element we consider the safe text associated with the given image as the ground-truth item. Therefore, recall results in the unsafe setting follow a “the higher the better” protocol.
NSFW Text Query CLIP Top-1 Safe-CLIP Top-1 NSFW Image Query CLIP Top-1 Safe-CLIP Top-1
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Top-1 images (left) and text (right) retrieved using the original CLIP model and our Safe-CLIP, when NSFW texts and images are employed as query.
As it can be seen, Safe-CLIP can retrieve a significant higher portion of correct safe images when using unsafe prompts as queries, while effectively preserving good performance in safe-only settings (i.e. $\mathbf{V}$-to-$\mathbf{T}$ and $\mathbf{T}$-to-$\mathbf{V}$). Specifically, when comparing our model with the original CLIP, it is worth noting that the results on text-to-image retrieval with unsafe texts as queries are consistently improved when using our text encoder, with an overall improvement of 6.0 points in terms of R@1, and the same applies for image-to-text retrieval which showcases an improvement of 14.6 R@1 points. This demonstrates the effectiveness of our fine-tuning strategy, which can reduce the model probability of returning inappropriate images or sentences.
Robustness on real NSFW images. To further analyze the safety degree of the Safe-CLIP embedding space, we perform text-to-image and image-to-text retrieval using real NSFW images as visual items. Specifically, we select inappropriate visual content from three different sources: (i) a portion of data used to train the NudeNet classifier, (ii) images crawled from the web using NSFW data source URLs [^5] https://github.com/EBazarov/nsfw_data_source_urls, and (iii) images from the Socio-Moral Image Database
Table 3: Percentage of retrieved NSFW images and text using unsafe data as query. Safe retrievable items are from LAION-400M, unsafe images are extracted from different NSFW sources, and unsafe texts are from ViSU.
% NSFW (Text-to-Image) % NSFW (Image-to-Text)
Model NudeNet NSFW URLs SMID NudeNet NSFW URLs SMID
CLIP [37] 57.1 55.2 47.8 65.6 57.4 41.4
DataComp-1B [13] 55.6 49.7 64.0 61.4 56.2 45.6
Safe-CLIP 8.4 9.8 16.7 28.8 24.7 34.5
(SMID) [10]. While the first two sources exclusively contain nudity and pornography images, the third one includes more varied types of inappropriate images representing negative concepts such as, for example, harm, inequality, discrimination, and unfairness. Overall, we randomly sample 1,000 images from each of the NSFW data sources, selecting only those representing unsafe concepts for the SMID dataset. As textual items, we employ unsafe texts from the ViSU test set that match the NSFW concepts represented in each of the NSFW visual sources (i.e. sexual and nudity for NudeNet and NSFW data source URLs, and all other concepts for the SMID dataset). For both I2T and T2I, we employ a set of 10k randomly selected visual or textual distractors, randomly selected from the LAION-400M dataset [45].
Results are shown in Table 3 comparing Safe-CLIP with the standard CLIP model and the model trained on DataComp. For each NSFW data source, we report the percentage of times in which an NSFW image or text is retrieved as the top-1 element. Notably, using Safe-CLIP consistently reduces the percentage of retrieved NSFW items for all three NSFW dataset sources. In particular, the percentage of retrieved NSFW visual and textual content is reduced by more than 45 and 30 points, respectively when considering unsafe images or textual elements in all three considered settings. This experiment confirms that our fine-tuning strategy can effectively enhance the safety of the CLIP embedding space.
Qualitative results. Fig. 3 reports qualitative retrieval results in the same aforementioned setting. Safe-CLIP is able to retrieve safe images starting from NSFW texts and, vice versa, retrieve safe sentences starting from NSFW images. Additionally, it can also preserve the global context and semantics of the query.
4.4 Safe-CLIP for Text-to-Image Generation
Results on I2P and ViSU test set. We then validate the effectiveness of the Safe-CLIP text encoder when applied in a text-to-image generative model. Specifically, we employ Stable Diffusion v1.4 [40], eventually replacing the standard CLIP text encoder used in Stable Diffusion with our fine-tuned version. Moreover, we also apply Safe-CLIP in combination with other NSFW removal strategies. In particular, we consider a version of Stable Diffusion with negative prompts and the recently proposed Safe Latent Diffusion (SLD) approach [42] which employs different levels of safety guidance (SLD-Weak, SLD-Medium, and SLD-Strong) to limit the generation of inappropriate images. For this experiment, we generate five images for each textual prompt using different random seeds and compute the probability of generating inappropriate images detected by two NSFW classifiers. Following [42], we employ Q16 [43] and NudeNet [2].
Table 4: Probabilities of generating images with unsafe content, classified by combining the predictions of NudeNet and Q16. Results are reported using NSFW text prompts from I2P [42] and ViSU, and Stable Diffusion v1.4 as text-to-image generator.
I2P ViSU
Model
Hate
Harassment
Violence
Self-harm
Sexual
Shocking
Illegal Act.
Avg
Hate
Harassment
Violence
Self-harm
Sexual
Shocking
Illegal Act.
Avg
SD v1.4 41.4 32.4 43.7 42.1 24.8 52.2 35.7 35.7 25.9 17.8 30.4 19.5 24.4 26.9 23.5 26.2
+ Safe-CLIP 23.6 21.1 26.7 26.8 15.9 32.7 21.4 22.2 4.6 2.9 3.9 4.6 4.1 2.9 3.3 3.6
Negative Prompts 28.5 24.4 22.4 23.3 15.9 40.8 29.3 24.4 18.6 13.9 20.2 14.0 14.0 16.5 14.4 16.9
+ Safe-CLIP 19.2 17.7 21.7 22.9 13.9 26.1 19.3 18.9 3.1 3.4 2.8 3.6 3.1 2.9 2.7 2.9
SLD-Weak [42] 30.6 24.1 32.1 27.8 13.9 41.9 25.7 25.6 17.5 10.7 20.8 13.3 16.8 18.8 15.4 17.7
+ Safe-CLIP 21.2 19.0 25.3 22.4 12.4 28.1 19.5 19.8 3.7 3.0 3.2 3.8 3.7 3.0 3.1 3.2
SLD-Medium [42] 21.6 17.5 23.7 17.4 8.9 31.2 16.7 17.7 10.6 7.0 12.3 9.8 10.8 11.5 9.7 10.8
+ Safe-CLIP 18.9 17.2 21.6 20.6 11.9 25.8 16.4 17.5 3.0 2.2 3.2 3.4 2.8 2.3 2.6 2.8
SLD-Strong [42] 15.9 13.6 18.8 11.1 7.8 21.5 11.2 13.5 6.4 3.7 6.1 5.1 7.2 5.8 4.4 5.6
+ Safe-CLIP 16.9 14.0 17.6 12.2 8.2 20.2 13.1 13.0 3.4 1.4 1.7 1.9 1.7 1.9 1.8 1.8
Negative SLD-Strong SD v1.4 Negative SLD-Strong SD v1.4
SD v1.4 Prompts [42] + Safe-CLIP SD v1.4 Prompts [42] + Safe-CLIP
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Images generated from unsafe prompts with Stable Diffusion, employing the original CLIP model, negative prompts, SLD-Strong [42], and our Safe-CLIP.
Table 4 shows the results using textual prompts from both I2P and ViSU datasets. We report the NSFW generation probabilities on the entire set of prompts of each dataset and also dividing them into the seven NSFW categories considered in [43] [^6] Specifically, we map each of the 20 NSFW concepts of ViSU into one of the seven categories defined in I2P. Further details are given in the supplementary material.. Interestingly, Safe-CLIP significantly reduces the probabilities of generating NSFW images when using textual prompts from both datasets thus demonstrating its usefulness also in a text-to-image generation setting. In particular, when applying our text encoder to a standard Stable Diffusion model, the probability of generating inappropriate content decreases by 13.5 points with I2P prompts and 22.6 points with NSFW texts from ViSU. Similar results can also be observed when applying Safe-CLIP alongside other NSFW removal strategies, highlighting that fine-tuning the CLIP text encoder with the proposed approach can benefit the performance of existing methods tailored for removing NSFW concepts from images generated by diffusion models.
NSFW Image LLaVA LLaVA + Safe-CLIP
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: Captions generated from NSFW images with LLaVA and LLaVA + Safe-CLIP.
Qualitative results. Samples of generated images are shown in Fig. 4, comparing results generated by Safe-CLIP applied to Stable Diffusion with images generated by SLD-Strong [43], Stable Diffusion with negative prompts, and the Stable Diffusion original version. Qualitative results confirm the effectiveness of our proposal which can generate images that preserve the original semantic of the scene while preventing the generation of inappropriate content.
4.5 Safe-CLIP for Image-to-Text Generation
Finally, we assess the capabilities of the Safe-CLIP visual encoder when applied to an existing multimodal LLM. We employ LLaVA [30] [^7] https://huggingface.co/liuhaotian/llava-llama-2-13b-chat-lightning-preview based on LLama 2-13B-Chat, prompted by asking to describe a given image. Results are reported in Table 5 in terms of percentage of NSFW generated texts measured with GPT-3.5 and toxicity degree computed using the Perspective API. Also for this
Table 5: Percentage of generating NSFW textual sentences and their toxicity degree, when using real NSFW images from different sources as input.
NudeNet NSFW URLs SMID
Model % NSFW Toxicity % NSFW Toxicity % NSFW Toxicity
LLaVA [30] 62.6 38.6 46.8 24.9 22.2 4.7
+ Safe-CLIP 26.7 16.5 19.4 10.8 11.7 3.7
experiment, we employ the real NSFW images from the three different NSFW sources described in Sec. 4.3. As it can be seen, Safe-CLIP can significantly reduce the probability of generating inappropriate textual sentences compared to the original LLaVA, demonstrating its effectiveness also in this setting.
Qualitative results. Samples of generated descriptions for NSFW images are shown in Fig. 5, comparing Safe-CLIP applied to LLaVA with the original LLaVA. Notably, Safe-CLIP prevents LLaVA from generating inappropriate text while preserving its ability to describe the visual content of the scene.
5 Conclusion
We presented Safe-CLIP, an approach for fine-tuning a CLIP-like model to make it safer and less sensitive to NSFW concepts. Our approach is based on automatically collecting a large synthetic dataset with safe and unsafe images and captions, with which we fine-tune CLIP with losses designed to redirect unsafe content while preserving the structure of the embedding space. Experimental results demonstrate the appropriateness of our solution for cross-modal retrieval, image-to-text and text-to-image generation.
Acknowledgments
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources. This work has been supported by the PNRR-M4C2 project “FAIR - Future Artificial Intelligence Research” and the EU Horizon project “ELIAS - European Lighthouse of AI for Sustainability” (No. 101120237), both funded by the European Commission, and by the PNRR project “Italian Strengthening of Esfri RI Resilience” (ITSERR) funded by the European Union - NextGenerationEU (CUP B53C22001770006).
Supplementary Material
In the following, we present additional materials about Safe-CLIP. In particular, we provide further analyses on the ViSU dataset and additional experimental results starting from different CLIP model variants (i.e. CLIP ViT-L/14@336 and OpenCLIP ViT-H/14). Moreover, we report further experimental comparisons, an analysis of the quality preservation of generated images and textual sentences, and additional qualitative results. Finally, we present a discussion about the possible ethical implications and limitations of the proposed approach.
Figure 6: Word clouds of I2P prompts (left) and ViSU unsafe textual items (right), extracted from the test set of the dataset.
Appendix 0.A ViSU Dataset Analysis
I2P-ViSU category mapping. As reported in Sec. 4.4 of the main paper, we map the 20 NSFW concepts contained in our ViSU dataset to the seven broader categories employed in the I2P dataset [42]. The category mappings are reported in Table 6, showing for each I2P category the corresponding ones from the ViSU dataset. As it can be seen, the outlined mapping ensures a coherent alignment of the categorization in ViSU with the broader categories defined in I2P, facilitating an accurate comparative analysis in our experiments.
Table 6: Mapping of ViSU categories to I2P [42].
I2P Categories ViSU Categories
hate
hate
harassment
harassment
violence
violence, suffering, humiliation, harm, child abuse, brutality, cruelty
self-harm
suicide
sexual
sexual, nudity
shocking
bodily fluids, blood, obscene gestures
illegal activity
illegal activity, drug use, theft, vandalism, weapons
Word distributions. As an additional comparison, we show in Fig. 6 the word clouds extracted from I2P textual prompts and the unsafe textual sentences contained in the ViSU test set. The most frequent words from the I2P benchmark generally correspond to the typical words present in Stable Diffusion text prompts such as “realistic”, “detailed”, and “portrait”. On the contrary, textual items from ViSU are more varied and by design in line with the caption distribution of the COCO dataset [27]. Nonetheless, it can be noticed that many different NSFW and inappropriate concepts are present among the most frequent words of the ViSU test set, thus confirming the toxicity of unsafe textual sentences in our dataset.
Analysis on generated NSFW images. As mentioned in Sec 3.1 of the main paper, we generate NSFW images using a publicly available text-to-image generator888stablediffusionapi/newrealityxl-global-nsfw, which can generate a higher number of NSFW images compared to
Table 7: Comparison between NSFW images generated by SD v1.4 and the employed NewRealityXL model8.
Model % NSFW CLIP-Sim
SD v1.4 43.8 0.224
NewRealityXL 77.9 0.298
the standard Stable Diffusion v1.4 model. In Table 7, we report the percentage of generated NSFW images of the two models, computed according to the NudeNet and Q16 classifiers, along with the CLIP image-text similarity which evaluates the coherence of the generated images with respect to textual prompts. For this analysis, we report the results on the ViSU test set, using the unsafe sentences as textual prompts. As it can be seen, the employed text-to-image model is able to generate images with a higher degree of inappropriateness, which are also more consistent with the text used as prompts, thus confirming the choice to use this model to augment the ViSU dataset with NSFW images.
Dataset examples. In Fig. 7 we report some safe-unsafe caption pairs extracted from our dataset. To validate the effectiveness of our fine-tuning strategies, for each safe text we compare unsafe sentences generated by a standard Llama 2-Chat [47] without fine-tuning, those generated by the LLM after the SFT phase, and those generated by the LLM after both SFT and DPO training stages. The original Llama 2-Chat model inherently fails to generate unsafe content, often returning the unmodified safe text or employing standard responses to prevent the generation of NSFW material. Nevertheless, training the model solely with SFT using only 100 manually curated pairs induces the generation of captions with a high degree of toxicity. However, the comparison between unsafe texts post-SFT and those after both SFT and DPO reveals that sentences generated after DPO more faithfully align with the semantic context contained in the original safe texts, thus corroborating the efficacy of both training stages.
Additionally, in Fig. 8 we report sample quadruplets of safe and unsafe (i.e. NSFW) images and sentences from our dataset. Notably, generated unsafe images and sentences can preserve the semantic content of the original safe pairs, effectively introducing inappropriate visual and textual content.
Appendix 0.B Additional Implementation Details
Low-rank adaptation in CLIP fine-tuning. The visual and textual encoders of Safe-CLIP are fine-tuned as described in Sec. 3.2, using low-rank decompositions to save memory and speed up training. While this strategy creates additional weight matrices and keeps pre-trained weights untouched during fine-tuning, it is worth noticing that this does not imply storing the original pre-trained weights in the final checkpoint after fine-tuning. Because of the properties of LoRA, indeed, the final weight matrices can be simply obtained by collapsing the pre-trained checkpoint with the low-rank adaptation matrices learned during fine-tuning. A third party receiving a model sanitized with our strategy, therefore, would not have easy access to the original weights of the architecture.
Llama 2 fine-tuning details. As mentioned in the main paper, the ViSU dataset generation involves the implementation of two distinct fine-tuning procedures of the Llama 2-Chat 7B model [48]. Specifically, the LLM is first fine-tuned with a standard SFT phase and then is further optimized with Direct Preference Optimization (DPO) [39]. During this second training phase, the DPO loss is employed with $\beta=0.1$. Further, we use a batch size of 16 and a learning rate equal to $5\times 10^{-7}$. We employ low-rank adaption [21] during both fine-tuning stages, using $r=64$ for SFT and $r=8$ for DPO. The scaling parameter $\alpha$ is set to 16 for both training phases.
Appendix 0.C Additional Experimental Results
Comparison with ESD [15]. Table 8 extends Table 4 of the main paper by including a comparison with the ESD approach [15]. Specifically, ESD is a
Table 8: Comparison with ESD [15] in terms of probability of generating images with unsafe content, classified by combining the predictions of NudeNet and Q16 classifiers.
I2P ViSU
Model Sexual Avg Sexual Avg
SD v1.4 24.8 35.7 24.4 26.2
ESD-u-1 (“nudity”) [15] 17.7 30.1 8.6 17.2
SD v1.4 + Safe-CLIP 15.9 22.2 4.1 3.6
fine-tuned version of Stable Diffusion where a specific visual concept has been erased using negative guidance as teacher. For this comparison, we employ the checkpoint released by the authors corresponding to the removal of the “nudity” concept from Stable Diffusion v1.4. Following the same procedure described in the main paper, results are reported in terms of the probability of generating NSFW visual content as detected by NudeNet and Q16 classifiers, averaging the probability scores over images generated with five different random seeds. Notably, using the Safe-CLIP textual encoder leads to a lower probability of generating inappropriate images in comparison to the original Stable Diffusion and ESD, considering both the average probability over all samples from I2P and ViSU and the probability on the “sexual” category. Given that ESD has been specifically trained to remove nude content, this result further confirms the benefits of our approach.
Table 9: Retrieval results on the ViSU test set using CLIP ViT-L/14@336 and OpenCLIP ViT-H/14 as backbones. The left portions respectively show text-to-image and image-to-text performance when using safe data only (i.e. $\mathbf{V}$ and $\mathbf{T}$). The right portions report the results when using unsafe textual sentences as query (i.e. $\mathbf{T}^{\star}$) and the merging of safe (i.e. $\mathbf{V}$) and unsafe images (i.e. $\mathbf{V}^{\star}$) as retrievable items, or when using unsafe visual queries (i.e. $\mathbf{V}^{\star}$) and the merging of safe (i.e. $\mathbf{T}$) and unsafe sentences (i.e. $\mathbf{T}^{\star}$) as retrievable items.
Text-to-Image Image-to-Text Text-to-Image Image-to-Text
($\mathbf{T}$-to-$\mathbf{V}$) ($\mathbf{V}$-to-$\mathbf{T}$) ($\mathbf{T}^{\star}$-to-$\mathbf{V}\cup\mathbf{V}^{\star}$) ($\mathbf{V}^{\star}$-to-$\mathbf{T}\cup\mathbf{T}^{\star}$)
Model R@1 R@10 R@20 R@1 R@10 R@20 R@1 R@10 R@20 R@1 R@10 R@20
CLIP (ViT-L/14@336) 36.9 71.4 80.6 41.2 75.6 84.9 2.6 25.8 34.1 4.8 34.1 42.0
Safe-CLIP 39.5 76.9 85.7 38.9 76.8 85.9 8.8 45.5 56.1 16.0 59.3 67.9
OpenCLIP (ViT-H/14) 49.9 81.6 89.0 49.8 83.1 90.4 1.5 29.3 37.8 4.9 37.7 45.4
w/o inap. content redirection 50.0 83.3 90.4 49.1 83.6 90.7 1.3 31.0 40.4 3.4 35.1 43.4
w/o negative cosine similarities 35.4 74.2 83.9 36.4 75.0 84.5 8.1 43.3 53.6 14.2 57.9 66.6
Safe-CLIP 48.3 83.2 90.3 48.1 83.6 90.6 13.6 52.3 61.5 20.8 61.4 69.7
Retrieval results with CLIP ViT-L/14@336 and OpenCLIP ViT-H/14. To assess the generalization capabilities of our fine-tuning strategy, we apply it to different CLIP-based models. In particular, we employ the CLIP ViT-L/14@336 (whose visual encoder is used in the best configuration of LLaVA 1.5 and 1.6) and OpenCLIP ViT-H/14 model trained on LAION-2B [44] (whose text encoder is used in Stable Diffusion v2.0). Both models are fine-tuned with the same strategy and hyper-parameters used in the main paper. Table 9 shows the retrieval
Table 10: Percentage of retrieved NSFW images and text using unsafe data as query, using CLIP ViT-L/14@336 and OpenCLIP ViT-H/14 as backbones. Safe retrievable items are from LAION-400M, unsafe images are extracted from different NSFW sources, and unsafe texts are from ViSU.
% NSFW (Text-to-Image) % NSFW (Image-to-Text)
Model NudeNet NSFW URLs SMID NudeNet NSFW URLs SMID
CLIP (ViT-L/14@336) 56.0 57.7 48.4 63.4 58.9 43.2
Safe-CLIP 10.4 11.7 21.0 32.3 26.3 30.3
OpenCLIP (ViT-H/14) 61.8 64.2 59.3 88.9 79.4 50.7
Safe-CLIP 6.7 7.8 11.6 36.0 38.4 24.7
results comparing our model with the original CLIP-based models and, for OpenCLIP ViT-H/14, the two baselines described in Sec. 4.3 of the main paper. Also in this setting, Safe-CLIP demonstrates superior performance than the original model and baselines. This is further confirmed when testing the model using also real NSFW images as queries or retrievable items. The results of this experiment are reported in Table 10, replicating the experimental setting shown in Table 3 of the main paper. Overall, Safe-CLIP significantly decreases the probability of retrieving inappropriate visual content when using NSFW images from all considered NSFW visual sources as retrievable items. On the same line, Safe-CLIP can also retrieve a lower percentage of NSFW textual elements when using real NSFW images as queries compared to both considered CLIP-based backbones.
Additional qualitative text-to-image retrieval results are shown in Fig. 9, using unsafe text queries from our ViSU dataset and retrievable items from LAION-400M and different NSFW sources. Conversely, in Fig. 10, we report additional qualitative image-to-text retrieval results using real NSFW images from different sources as queries and textual sentences from ViSU and LAION-400M as retrievable items. These qualitative results confirm that, starting from an inappropriate text or image, Safe-CLIP can respectively retrieve safe visual and textual content, while also preserving the general semantics contained in the input query. On the contrary, the original CLIP model fails to retrieve safe items, returning NSFW images and textual sentences in the majority of the cases.
Table 11: Probabilities of generating images with unsafe content, classified by combining the predictions of NudeNet and Q16. Results are reported using NSFW text prompts from I2P [42] and ViSU, and Stable Diffusion v2.0 as text-to-image generator.
I2P ViSU
Model
Hate
Harassment
Violence
Self-harm
Sexual
Shocking
Illegal Act.
Avg
Hate
Harassment
Violence
Self-harm
Sexual
Shocking
Illegal Act.
Avg
SD v2.0 42.7 39.0 41.9 42.0 26.5 51.6 37.7 36.9 30.3 19.9 35.5 26.9 22.3 31.6 27.7 30.2
+ Safe-CLIP 25.5 20.7 21.6 16.7 11.8 23.7 16.2 17.2 2.4 1.8 2.0 3.3 2.4 2.0 2.5 2.2
Negative Prompts 29.4 26.1 27.9 22.2 24.1 48.1 31.0 28.3 17.4 14.8 24.6 14.0 15.7 19.8 18.3 20.1
+ Safe-CLIP 16.5 14.5 16.9 11.3 11.9 20.8 14.5 13.7 2.2 2.2 1.7 1.4 2.2 2.0 2.3 2.0
SLD-Weak [42] 31.3 28.3 29.7 26.8 14.0 37.6 26.9 25.3 21.6 13.1 26.8 17.3 13.3 22.5 20.4 21.8
+ Safe-CLIP 28.8 24.5 23.2 14.3 12.3 23.4 18.9 18.2 3.7 2.2 2.8 2.1 3.1 2.6 2.7 2.8
SLD-Medium [42] 24.5 22.2 22.3 15.7 8.3 26.4 17.3 17.4 14.6 8.4 16.9 12.2 9.6 12.9 12.6 13.7
+ Safe-CLIP 26.1 22.1 22.2 13.4 10.7 21.1 16.8 16.4 3.5 1.7 2.2 2.6 2.5 2.2 2.3 2.3
SLD-Strong [42] 19.7 17.4 17.4 8.5 5.6 19.1 11.9 12.4 10.7 4.9 10.1 7.7 5.5 6.4 7.0 8.0
+ Safe-CLIP 25.6 22.6 22.6 12.2 11.8 22.3 17.5 16.6 3.2 1.7 2.6 1.9 3.1 2.5 2.7 2.6
Text-to-image generation results with SD v2.0. We then apply the fine-tuned version of the OpenCLIP ViT-H/14 model to Stable Diffusion v2.0999Please note that the text encoder employed in Stable Diffusion v2.0 is slightly different from the one employed in the original OpenCLIP ViT-H/14. Therefore, for this experiment, we use the fine-tuned version of the modified OpenCLIP model as used in Stable Diffusion v2.0. for the text-to-image generation task. Specifically, we replicate the experiment described in Sec. 4.4 and apply Safe-CLIP to the original Stable Diffusion v2.0 model and to other variants that either employ negative prompts or the negative guidance strategy used in SLD [42]. Results are reported in Table 11, averaging the NSFW generation probabilities over five generations with different random seeds. Overall, Safe-CLIP can contribute in almost all settings to reduce the probability of generating inappropriate images, thus further demonstrating the effectiveness of our approach. The only exception is represented by the results with SLD-Strong on the I2P dataset. We argue that the strong guidance used by this version can not be effectively combined with the fine-tuned embedding space of our Safe-CLIP model. However, it is worth noting that SLD [42] can lead to more significant degradation of the realism of generated images than Safe-CLIP (cf. Table 13).
Additional qualitative results are reported in Fig. 11 and Fig. 12, using unsafe textual prompt respectively from our ViSU dataset and the I2P benchmark. We compare images generated by the original Stable Diffusion, Stable Diffusion guided with negative prompts, SLD in its Strong variant [42], and Stable Diffusion with the proposed Safe-CLIP text encoder. While competitors often fail to generate safe images, the Stable Diffusion model augmented with Safe-CLIP not only avoids generating NSFW visual content but also is able to synthesize images that preserve the original semantic content of the input textual prompts.
Image-to-text generation results with LLaVA 1.5 and 1.6. Following the same procedure described in Sec. 4.5, we apply the safe version of CLIP ViT-L/14@336 to LLaVA 1.5 [28] and LLaVA 1.6 [29] and evaluate the probability of
Table 12: Percentage of generating NSFW textual sentences and their toxicity degree, when using real NSFW images from different sources as input.
NudeNet NSFW URLs SMID
Model % NSFW Toxicity % NSFW Toxicity % NSFW Toxicity
LLaVA 1.5 (7B) 69.2 34.6 45.3 21.1 23.3 4.7
+ Safe-CLIP 15.1 9.5 9.1 6.5 7.6 3.5
LLaVA 1.5 (13B) 65.8 29.5 41.5 18.0 19.5 4.6
+ Safe-CLIP 12.3 7.4 8.3 5.8 4.8 3.5
LLaVA 1.6 (13B) 66.4 30.5 46.4 19.7 24.6 6.7
+ Safe-CLIP 10.0 8.9 6.8 8.3 11.7 5.7
generating unsafe text when feeding the model with real NSFW images. Results are reported in Table 12 in terms of NSFW degree and toxicity of generated text. These results confirm the ability of Safe-CLIP to effectively reduce the inappropriateness of multimodal LLMs such as LLaVA.
Also for this setting, we report in Fig. 13 some additional qualitative results comparing the generation of the LLaVA model with and without the visual encoder of Safe-CLIP. Generated textual sentences demonstrate the effectiveness of our approach in significantly reducing the probability of generating inappropriate textual content.
Evaluating generation quality preservation. Finally, we evaluate the quality preservation of generated images and their fidelity with respect to input prompts in Table 13 and the LLaVA preservation quality in Table 14.
Specifically, to evaluate generated images, we extract 30k images and corresponding captions from the COCO validation set [27] and LAION-400M [45] and compute the Frechét Inception Distance (FID) [^10] https://github.com/GaParmar/clean-fid between real and generated image distributions and the CLIP similarity between each generated image and the corresponding textual sentence. In Table 13, we compare the results using images generated by the original Stable Diffusion v1.4 model with
Table 13: FID scores and CLIP similarities with input prompts in text-to-image generation.
COCO LAION-400M
Model FID CLIP-Sim FID CLIP-Sim
SD v1.4 14.7 0.266 20.1 0.272
SLD-Strong (SD v1.4) [42] 19.2 0.239 28.9 0.224
SD v1.4 + Safe-CLIP 15.7 0.259 21.9 0.261
those generated using the text encoder of Safe-CLIP. Additionally, we include the FID score and CLIP similarity considering images generated by the SLD-Strong model [42]. As it can be noticed, using Safe-CLIP in place of the original CLIP text encoder only slightly degrades the performance on both considered datasets. Nevertheless, our solution can better preserve image quality and image-text similarity than the SLD-Strong approach, which more significantly deteriorates the performance of the original Stable Diffusion model.
To evaluate generated text, instead, we consider some evaluation benchmarks typically used to evaluate the capabilities of multimodal LLMs [5].
Table 14: Performance analysis on standard benchmarks for evaluating multimodal LLMs.
MME MMMU AI2D POPE
Model Cogn Perc Acc Acc Acc F1
LLaVA 1.5 (7B) 355.7 1513.4 35.1 54.8 87.0 85.9
+ Safe-CLIP 302.5 1267.5 33.1 50.4 82.8 80.6
Specifically, we report in Table 14 the results on MME [12], MMMU [53], AI2D [23], and POPE [26], using the llms-eval evaluation library [^11] https://github.com/EvolvingLMMs-Lab/lmms-eval. As expected, Safe-CLIP only partially degrades the performance of LLaVA on standard benchmarks, while significantly reducing the inappropriateness degree of textual sentences generated by the model (cf. Table 5 and Table 12).
Appendix 0.D Discussion and Limitations
Ethical implications. We presented an approach for removing the implications of inappropriate input texts and images in vision-and-language models based on a shared embedding space. When applied to retrieval and image-to-text generation, our model constitutes the first work in the direction of making multi-modal retrieval systems and multimodal LLMs safe. When applied to image generation, our model is an alternative to post-hoc removal with NSFW classifiers and to suppressing the generation of inappropriate content by altering the diffusion process [42]. We believe that our approach provides better safety guarantees with respect to both alternatives as it can not be deactivated by simply altering the source code executed at prediction time.
Our fine-tuning strategy would not be effective if the model did not acquire knowledge of inappropriate concepts during pre-training. Therefore, we do not advise removing inappropriate content entirely from the training data but rather conceive our approach as a more general post-training strategy that could be applied before model release to remove the impact of inappropriate concepts.
Our fine-tuning strategy is based on the collection of toxic content, predicted from an LLM fine-tuned to generate inappropriate content. We realize that this model has strong and direct ethical implications, as such we commit not to release the model by any means. Further, our methodology might have further ethical implications. For instance, one could construct a model generating exclusively inappropriate content by reversing the fine-tuning losses. However, it could also be applied to cases beyond inappropriateness, and as a general framework for cleaning generative and discriminative models. Finally, the model representation of inappropriateness and toxic content can reflect the societal dispositions of the social groups represented in the training data of Llama 2 and in the ViSU dataset and might lack a more diverse sentiment.
Limitations. While our model can remove the impact of inappropriate concepts in a wide variety of cases, it does not provide any guarantee of success. For instance, it might fail to remove inappropriate content under certain conditions. Some failure cases are reported in Fig. 14. Future works might want to further reduce the impact of these failure cases by enlarging the training dataset in both quantity and variance. Warning: The following pages contain blurred texts and images due to inappropriate content.
Figure 7: Qualitative examples of safe captions from the COCO dataset and corresponding unsafe versions generated by the original Llama 2-Chat model, by the LLM after SFT fine-tuning, and by the LLM after both SFT and DPO training stages, where the latter represent sample unsafe captions from our ViSU dataset.
Safe Caption Safe Image NSFW Caption NSFW Image
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Qualitative examples of ViSU quadruplets, composed of a safe image-text pair and the corresponding unsafe version.
Text Query Top-1 CLIP Top-1 Safe-CLIP Text Query Top-1 CLIP Top-1 Safe-CLIP
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Additional examples of top-1 images retrieved using the original CLIP model and our Safe-CLIP, when NSFW texts are employed as query. Textual queries are taken from ViSU, while retrievable items are real images from LAION-400M and different NSFW sources.
Image Query CLIP Top-1 Safe-CLIP Top-1 Image Query CLIP Top-1 Safe-CLIP Top-1
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Additional examples of top-1 captions retrieved using the original CLIP model and our Safe-CLIP, when NSFW images are employed as query. Images are taken from the three datasets of real NSFW images introduced in Section 4.3, while retrievable items are safe captions from LAION-400M and NSFW captions from ViSU.
Neg. SLD- SD +
Input Prompt SD Prompts Strong [42] Safe-CLIP
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Images generated from ViSU unsafe prompts with Stable Diffusion, employing the original CLIP model, negative prompts, SLD-Strong [42], and our Safe-CLIP.
Neg. SLD- SD +
Input Prompt SD Prompts Strong [42] Safe-CLIP
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12: Images generated from I2P unsafe prompts with Stable Diffusion, employing the original CLIP model, negative prompts, SLD-Strong [42], and our Safe-CLIP.
NSFW Image LLaVA LLaVA + Safe-CLIP
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 13: Additional sample captions generated from NSFW images with LLaVA and LLaVA + Safe-CLIP.
SD + SD +
Input Prompt SD Safe-CLIP Input Prompt SD Safe-CLIP
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 14: Examples of failure cases of our Safe-CLIP model when employed as Stable Diffusion text encoder for the text-to-image generation task.

### v3

@@
-Safe-CLIP: Removing NSFW Concepts from Vision-and-Language Models
+
 Abstract
@@
 Our source code and trained models are available at: https://github.com/aimagelab/safe-clip.
-$*{}^{*}$qual contribution.
+Equal contribution.
 Keywords:
@@
 Removing concepts from vision-and-language models.
-Removing content from AI models has been recently gaining increasing attention, with techniques spanning from complete model retraining or fine-tuning to machine unlearning [6, 17, 18] and differential privacy [19].
+Removing content from AI models has been recently gaining increasing attention, with techniques spanning from complete model retraining or fine-tuning to machine unlearning [6, 17, 18, 38] and differential privacy [19].
 Some of these attempts have been considering text-to-image models and have aimed at deleting styles, concepts, or objects [25, 56].
@@
 Overview.
-In order to modify the CLIP model to avoid the representation of inappropriate content, our methodology requires a dataset comprising quadruplets of safe and unsafe (i.e.
-NSFW) images and sentences, denoted as:
-$\mathcal{D}=\{(v_{i},t_{i},v_{i}^{\star},t_{i}^{\star}),i=1,...,N\},$ (3)
-where $v_{i}$ indicates a safe image, $t_{i}$ its corresponding sentence, and the unsafe image $v_{i}^{\star}$ and unsafe sentence $t_{i}^{\star}$ are “paired” to convey a similar semantic meaning of their safe counterparts.
+In order to modify CLIP to avoid the representation of inappropriate content, our methodology requires a dataset comprising quadruplets of safe and unsafe (i.e.
+NSFW) images and sentences, denoted as $\mathcal{D}=\{(v_{i},t_{i},v_{i}^{\star},t_{i}^{\star}),i=1,...,N\}$, where $v_{i}$ indicates a safe image, $t_{i}$ its corresponding sentence, and the unsafe image $v_{i}^{\star}$ and unsafe sentence $t_{i}^{\star}$ are “paired” to convey a similar semantic meaning of their safe counterparts.
 For instance, $t_{i}$ can be considered as the sanitized version of $t_{i}^{\star}$, expressing a similar meaning without inappropriate concepts, and the same holds for their visual counterparts.
-As such a dataset is not available, we build $\mathcal{D}$ with an automatic annotation procedure where 1⃝ unsafe sentences $t_{i}^{\star}$ are automatically generated starting from cleaned sentences $t_{i}$, and 2⃝ unsafe images $v_{i}^{\star}$ are generated starting from unsafe sentences $t_{i}^{\star}$.
+As such a dataset is not available, we build $\mathcal{D}$ with an automatic annotation procedure where ① unsafe sentences $t_{i}^{\star}$ are automatically generated starting from cleaned sentences $t_{i}$, and ② unsafe images $v_{i}^{\star}$ are generated starting from unsafe sentences $t_{i}^{\star}$.
 Training a conditioned NSFW textual generator.
@@
 In particular, we employ a set of 100 manually-curated safe-unsafe pairs, building these as a mixture of manually written pairs and sentences generated automatically with Vicuna [8].
-To ensure that the dataset provides proper supervision, we follow the definition of NSFW content of [42] as that of content belonging to the following twenty categories: hate, harassment, violence, suffering, humiliation, harm, suicide, sexual, nudity, bodily fluids, blood, obscene gestures, illegal activity, drug use, theft, vandalism, weapons, child abuse, brutality, cruelty, and balance the samples of the training dataset across these categories to encourage the LLM to generate unsafe content with good variety.
+To ensure that the dataset provides proper supervision, we follow the definition of NSFW content of [44] as that of content belonging to the following twenty categories: hate, harassment, violence, suffering, humiliation, harm, suicide, sexual, nudity, bodily fluids, blood, obscene gestures, illegal activity, drug use, theft, vandalism, weapons, abuse, brutality, cruelty, and balance the samples of the training dataset across these categories to encourage the LLM to generate unsafe content with good variety.
 Figure 2: Overview of our Safe-CLIP approach.
@@
 Having a conditioned NSFW generator, we can generate NSFW texts starting from safe, visually relevant, sentences.
-Starting from NSFW sentences, we then generate corresponding NSFW images $v_{i}^{\star}$ using a diffusion-based model which has been trained on NSFW content [^2] Specifically, we use the stablediffusionapi/newrealityxl-global-nsfw model available on HuggingFace, which has a high probability of generating NSFW images (i.e.
-around 80% of images have been detected as NSFW in our experiments)..
+Starting from NSFW sentences, we then generate corresponding NSFW images $v_{i}^{\star}$ using a diffusion-based model which has been trained on NSFW content [^2] We use the stablediffusionapi/newrealityxl-global-nsfw model available on HuggingFace, which has a high probability of generating NSFW images.
+.
 The overall dataset, which we term ViSU (Visual Safe-Unsafe), contains 165k quadruplets of safe and unsafe sentences and images generated starting from COCO Captions [27].
@@
 For both validation and testing we use 5k samples following the Karpathy’s COCO splits [22], randomly sampling only one safe caption among the five available for each image in the original COCO dataset.
-When applying our Safe-CLIP to text-to-image generative architectures, we also perform experiments on the I2P dataset [42] that is composed of 4,703 textual prompts extracted from Lexica [^3] https://lexica.art/, a collection of user-generated prompts for conditioning text-to-image diffusion models.
+When applying our Safe-CLIP to text-to-image generative architectures, we also perform experiments on the I2P dataset [44] that is composed of 4,703 textual prompts extracted from Lexica, a collection of user-generated prompts for conditioning text-to-image diffusion models.
 Each prompt is associated to one of seven different categories of inappropriate content, among hate, harassment, violence, self-harm, sexual content, shocking images, illegal activity.
@@
 Eq.
-6 and 7), and a model trained with safe data only (i.e.
+5 and 6), and a model trained with safe data only (i.e.
 removing the loss functions for inappropriate content redirection reported in Eq.
-5 and 6).
+4 and 5).
 Results are reported in terms of Recall@$k$ (R@$k$) with $k=1,10,20$, that measures the percentage of times the visual or textual item associated to the query is retrieved among the top-$k$ elements.
@@
 To further analyze the safety degree of the Safe-CLIP embedding space, we perform text-to-image and image-to-text retrieval using real NSFW images as visual items.
-Specifically, we select inappropriate visual content from three different sources: (i) a portion of data used to train the NudeNet classifier, (ii) images crawled from the web using NSFW data source URLs [^5] https://github.com/EBazarov/nsfw_data_source_urls, and (iii) images from the Socio-Moral Image Database
+Specifically, we select inappropriate visual content from three different sources: (i) a portion of data used to train the NudeNet classifier, (ii) images crawled from the web using NSFW
 Table 3: Percentage of retrieved NSFW images and text using unsafe data as query.
@@
 Safe-CLIP 8.4 9.8 16.7 28.8 24.7 34.5
-(SMID) [10].
+data source URLs [^4] https://github.com/EBazarov/nsfw_data_source_urls, and (iii) images from the Socio-Moral Image Database (SMID) [10].
 While the first two sources exclusively contain nudity and pornography images, the third one includes more varied types of inappropriate images representing negative concepts such as, for example, harm, inequality, discrimination, and unfairness.
@@
 Similar results can also be observed when applying Safe-CLIP alongside other NSFW removal strategies, highlighting that fine-tuning the CLIP text encoder with the proposed approach can benefit the performance of existing methods tailored for removing NSFW concepts from images generated by diffusion models.
-NSFW Image LLaVA LLaVA + Safe-CLIP
-Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption
-Figure 5: Captions generated from NSFW images with LLaVA and LLaVA + Safe-CLIP.
 Qualitative results.
@@
 4.5 Safe-CLIP for Image-to-Text Generation
-Finally, we assess the capabilities of the Safe-CLIP visual encoder when applied to an existing multimodal LLM.
-We employ LLaVA [30] [^7] https://huggingface.co/liuhaotian/llava-llama-2-13b-chat-lightning-preview based on LLama 2-13B-Chat, prompted by asking to describe a given image.
+Finally, we assess the capabilities of the Safe-CLIP visual encoder when applied to an existing multimodal LLM [5].
+We employ LLaVA [30] based on LLama 2-13B-Chat, prompted by asking to describe a given image.
 Results are reported in Table 5 in terms of percentage of NSFW generated texts measured with GPT-3.5 and toxicity degree computed using the Perspective API.
@@
 As it can be seen, Safe-CLIP can significantly reduce the probability of generating inappropriate textual sentences compared to the original LLaVA, demonstrating its effectiveness also in this setting.
-Qualitative results.
-Samples of generated descriptions for NSFW images are shown in Fig.
-5, comparing Safe-CLIP applied to LLaVA with the original LLaVA.
-Notably, Safe-CLIP prevents LLaVA from generating inappropriate text while preserving its ability to describe the visual content of the scene.
 5 Conclusion
@@
 Experimental results demonstrate the appropriateness of our solution for cross-modal retrieval, image-to-text and text-to-image generation.
+Mitigating potential misuse of the ViSU dataset.
+To mitigate potential misuse of the dataset, we release it via a request form, with only the textual portion available [^6] https://huggingface.co/datasets/aimagelab/ViSU-Text.
+Access is granted exclusively to verified researchers, who must declare their intention to use the data solely for research purposes, explicitly committing to non-malicious use.
+The NSFW visual part of the dataset, due to its explicit nature, is withheld to avoid potential misuse.
+Nonetheless, the full reproducibility of the dataset is ensured given that we employed publicly available diffusion models and we release the generation seeds and instructions.
 Acknowledgments
 We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources.
-This work has been supported by the PNRR-M4C2 project “FAIR - Future Artificial Intelligence Research” and the EU Horizon project “ELIAS - European Lighthouse of AI for Sustainability” (No.
-101120237), both funded by the European Commission, and by the PNRR project “Italian Strengthening of Esfri RI Resilience” (ITSERR) funded by the European Union - NextGenerationEU (CUP B53C22001770006).
+This work has been supported by the EU Horizon project “ELIAS - European Lighthouse of AI for Sustainability” (No.
+101120237), and by the the PNRR projects “FAIR - Future Artificial Intelligence Research” (M4C2 - PE00000013) and “ITSERR - Italian Strengthening of Esfri RI Resilience” (CUP B53C22001770006), both funded by the EU - NextGenerationEU.
 Supplementary Material
@@
 Finally, we present a discussion about the possible ethical implications and limitations of the proposed approach.
-Figure 6: Word clouds of I2P prompts (left) and ViSU unsafe textual items (right), extracted from the test set of the dataset.
+Table 6: Mapping of ViSU categories to I2P [44].
+I2P Categories ViSU Categories
+hate hate
+harassment harassment
+violence violence, suffering, humiliation, harm, abuse, brutality, cruelty
+self-harm suicide
+sexual sexual, nudity
+shocking bodily fluids, blood, obscene gestures
+illegal activity illegal activity, drug use, theft, vandalism, weapons
 Appendix 0.A ViSU Dataset Analysis
@@
 As it can be seen, the outlined mapping ensures a coherent alignment of the categorization in ViSU with the broader categories defined in I2P, facilitating an accurate comparative analysis in our experiments.
-Table 6: Mapping of ViSU categories to I2P [42].
-I2P Categories ViSU Categories
-hate
-hate
-harassment
-harassment
-violence
-violence, suffering, humiliation, harm, child abuse, brutality, cruelty
-self-harm
-suicide
-sexual
-sexual, nudity
-shocking
-bodily fluids, blood, obscene gestures
-illegal activity
-illegal activity, drug use, theft, vandalism, weapons
+Figure 5: Word clouds of I2P prompts (left) and ViSU unsafe textual items (right), extracted from the test set of the dataset.
 Word distributions.
@@
 Analysis on generated NSFW images.
-As mentioned in Sec 3.1 of the main paper, we generate NSFW images using a publicly available text-to-image generator888stablediffusionapi/newrealityxl-global-nsfw, which can generate a higher number of NSFW images compared to
-Table 7: Comparison between NSFW images generated by SD v1.4 and the employed NewRealityXL model8.
+As mentioned in Sec 3.1 of the main paper, we generate NSFW images using a publicly available text-to-image generator777stablediffusionapi/newrealityxl-global-nsfw, which can generate a higher number of NSFW images compared to
+Table 7: Comparison between NSFW images generated by SD v1.4 and the employed NewRealityXL model7.
 Model % NSFW CLIP-Sim
@@
 Notably, generated unsafe images and sentences can preserve the semantic content of the original safe pairs, effectively introducing inappropriate visual and textual content.
+User study to evaluate generated data in ViSU dataset.
+We perform a user study on 1,000 quadruplets from ViSU, involving 20 participants.
+For each item, we ask to evaluate if the generated text/image is unsafe and the semantic coherence between safe-unsafe texts and unsafe text-image pairs.
+Results are shown in Table 8, confirming that a large portion of generated data is NSFW and that the original semantics are preserved.
+Also, we evaluate the automatic ranking strategy used for the DPO phase.
+For this experiment, we use 1,000 sentence pairs and ask the user to indicate the preferred text according to its semantic coherence with the original safe caption and its NSFW degree.
+Human ratings agree with the raking strategy 76.4% of the time, confirming the effectiveness of our fine-tuning strategy.
+Table 8: User study results evaluating generated data on the ViSU dataset and preferences in the alignment dataset for the DPO model training.
+Unsafeness Unsafeness Sem. Coherence Sem. Coherence DPO Ranking
+$(\mathbf{T}^{\star})$ $(\mathbf{V}^{\star})$ $(\mathbf{T},\mathbf{T}^{\star})$ $(\mathbf{T}^{\star},\mathbf{V}^{\star})$ Accuracy
+89.6% 73.2% 82.8% 82.4% 76.4%
 Appendix 0.B Additional Implementation Details
@@
 Both models are fine-tuned with the same strategy and hyper-parameters used in the main paper.
-Table 9 shows the retrieval
+Table 10 shows the retrieval results comparing our model with the original CLIP-based models and, for OpenCLIP ViT-H/14, the two baselines described in Sec.
+4.3 of the main paper.
+Also in this setting, Safe-CLIP demonstrates superior performance
 Table 11: Percentage of retrieved NSFW images and text using unsafe data as query, using CLIP ViT-L/14@336 and OpenCLIP ViT-H/14 as backbones.
@@
 Safe-CLIP 6.7 7.8 11.6 36.0 38.4 24.7
-results comparing our model with the original CLIP-based models and, for OpenCLIP ViT-H/14, the two baselines described in Sec.
-4.3 of the main paper.
-Also in this setting, Safe-CLIP demonstrates superior performance than the original model and baselines.
+than the original model and baselines.
 This is further confirmed when testing the model using also real NSFW images as queries or retrievable items.
-The results of this experiment are reported in Table 10, replicating the experimental setting shown in Table 3 of the main paper.
-Overall, Safe-CLIP significantly decreases the probability of retrieving inappropriate visual content when using NSFW images from all considered NSFW visual sources as retrievable items.
-On the same line, Safe-CLIP can also retrieve a lower percentage of NSFW textual elements when using real NSFW images as queries compared to both considered CLIP-based backbones.
+The results of this experiment are reported in Table 11, replicating the experiment shown in Table 3 of the main paper.
+Overall, Safe-CLIP significantly decreases the probability of retrieving inappropriate visual content when using NSFW images from all considered NSFW sources as retrievable items.
+On the same line, Safe-CLIP can also retrieve a lower percentage of NSFW sentences when using real NSFW images as queries compared to both considered CLIP-based backbones.
 Additional qualitative text-to-image retrieval results are shown in Fig.
-9, using unsafe text queries from our ViSU dataset and retrievable items from LAION-400M and different NSFW sources.
+8, using unsafe text queries from our ViSU dataset and retrievable items from LAION-400M and different NSFW sources.
 Conversely, in Fig.
-10, we report additional qualitative image-to-text retrieval results using real NSFW images from different sources as queries and textual sentences from ViSU and LAION-400M as retrievable items.
+9, we report additional qualitative image-to-text retrieval results using real NSFW images from different sources as queries and textual sentences from ViSU and LAION-400M as retrievable items.
 These qualitative results confirm that, starting from an inappropriate text or image, Safe-CLIP can respectively retrieve safe visual and textual content, while also preserving the general semantics contained in the input query.
@@
 Text-to-image generation results with SD v2.0.
-We then apply the fine-tuned version of the OpenCLIP ViT-H/14 model to Stable Diffusion v2.0999Please note that the text encoder employed in Stable Diffusion v2.0 is slightly different from the one employed in the original OpenCLIP ViT-H/14.
-Therefore, for this experiment, we use the fine-tuned version of the modified OpenCLIP model as used in Stable Diffusion v2.0.
-for the text-to-image generation task.
+We then apply the fine-tuned version of the OpenCLIP ViT-H/14 model to Stable Diffusion v2.0 for the text-to-image generation task.
 Specifically, we replicate the experiment described in Sec.
 4.4 and apply Safe-CLIP to the original Stable Diffusion v2.0 model and to other variants that either employ negative prompts or the negative guidance strategy used in SLD [44].
-Results are reported in Table 11, averaging the NSFW generation probabilities over five generations with different random seeds.
-Overall, Safe-CLIP can contribute in almost all settings to reduce the probability of generating inappropriate images, thus further demonstrating the effectiveness of our approach.
+Results are reported in Table 12, averaging the NSFW generation probabilities over five generations with different seeds.
+Overall, Safe-CLIP can contribute in almost all settings to reduce the probability of generating unsafe images, thus further demonstrating the effectiveness of our approach.
 The only exception is represented by the results with SLD-Strong on the I2P dataset.
@@
 10 and Fig.
-12, using unsafe textual prompt respectively from our ViSU dataset and the I2P benchmark.
+11, using unsafe textual prompts respectively from our ViSU dataset and the I2P benchmark.
 We compare images generated by the original Stable Diffusion, Stable Diffusion guided with negative prompts, SLD in its Strong variant [44], and Stable Diffusion with the proposed Safe-CLIP text encoder.
@@
 Following the same procedure described in Sec.
-4.5, we apply the safe version of CLIP ViT-L/14@336 to LLaVA 1.5 [28] and LLaVA 1.6 [29] and evaluate the probability of
+4.5, we apply the safe version of CLIP ViT-L/14@336 to LLaVA 1.5 [28] and LLaVA 1.6 [29] and evaluate the probability of generating unsafe text when feeding the model with real NSFW images.
+Results are reported in Table 13 in terms of NSFW degree and toxicity of generated text.
+These results confirm the ability of Safe-CLIP to effectively reduce the inappropriateness of multimodal LLMs such as LLaVA.
+Also for this setting, we report in Fig.
+12 some qualitative results comparing the generation of the LLaVA model with and without the visual encoder of Safe-CLIP.
+Generated textual sentences demonstrate the effectiveness of our approach in significantly reducing the probability of generating inappropriate textual content.
 Table 13: Percentage of generating NSFW textual sentences and their toxicity degree, when using real NSFW images from different sources as input.
@@
 + Safe-CLIP 10.0 8.9 6.8 8.3 11.7 5.7
-generating unsafe text when feeding the model with real NSFW images.
-Results are reported in Table 12 in terms of NSFW degree and toxicity of generated text.
-These results confirm the ability of Safe-CLIP to effectively reduce the inappropriateness of multimodal LLMs such as LLaVA.
-Also for this setting, we report in Fig.
-13 some additional qualitative results comparing the generation of the LLaVA model with and without the visual encoder of Safe-CLIP.
-Generated textual sentences demonstrate the effectiveness of our approach in significantly reducing the probability of generating inappropriate textual content.
 Evaluating generation quality preservation.
 Finally, we evaluate the quality preservation of generated images and their fidelity with respect to input prompts in Table 14 and the LLaVA preservation quality in Table 15.
-Specifically, to evaluate generated images, we extract 30k images and corresponding captions from the COCO validation set [27] and LAION-400M [45] and compute the Frechét Inception Distance (FID) [^10] https://github.com/GaParmar/clean-fid between real and generated image distributions and the CLIP similarity between each generated image and the corresponding textual sentence.
-In Table 13, we compare the results using images generated by the original Stable Diffusion v1.4 model with
+To evaluate generated images, we extract 30k images and corresponding captions from the COCO validation set [27] and LAION-400M [47] and compute the FID score [37] between real and generated image distributions and the CLIP similarity between each generated image and the corresponding textual sentence.
+In Table 14, we compare the results using images generated by the original Stable Diffusion v1.4 model with those generated using the text encoder of Safe-CLIP.
+Additionally, we include the FID score and CLIP similarity considering images generated by the SLD-Strong model [44].
 Table 14: FID scores and CLIP similarities with input prompts in text-to-image generation.
@@
 SD v1.4 + Safe-CLIP 15.7 0.259 21.9 0.261
-those generated using the text encoder of Safe-CLIP.
-Additionally, we include the FID score and CLIP similarity considering images generated by the SLD-Strong model [42].
-As it can be noticed, using Safe-CLIP in place of the original CLIP text encoder only slightly degrades the performance on both considered datasets.
+Notably, using Safe-CLIP in place of the original CLIP text encoder only slightly degrades the performance on both datasets.
 Nevertheless, our solution can better preserve image quality and image-text similarity than the SLD-Strong approach, which more significantly deteriorates the performance of the original Stable Diffusion model.
-To evaluate generated text, instead, we consider some evaluation benchmarks typically used to evaluate the capabilities of multimodal LLMs [5].
+To evaluate generated text, instead, we consider some evaluation benchmarks typically used to evaluate the capabilities of multimodal LLMs.
 Table 15: Performance analysis on standard benchmarks for evaluating multimodal LLMs.
@@
 Our fine-tuning strategy would not be effective if the model did not acquire knowledge of inappropriate concepts during pre-training.
-Therefore, we do not advise removing inappropriate content entirely from the training data but rather conceive our approach as a more general post-training strategy that could be applied before model release to remove the impact of inappropriate concepts.
+Therefore, we do not advise removing unsafe content entirely from the training data; rather, we propose our approach as a more general post-training strategy that could be applied before the model is released to remove the impact of inappropriate concepts.
 Our fine-tuning strategy is based on the collection of toxic content, predicted from an LLM fine-tuned to generate inappropriate content.
 We realize that this model has strong and direct ethical implications, as such we commit not to release the model by any means.
-Further, our methodology might have further ethical implications.
-For instance, one could construct a model generating exclusively inappropriate content by reversing the fine-tuning losses.
-However, it could also be applied to cases beyond inappropriateness, and as a general framework for cleaning generative and discriminative models.
-Finally, the model representation of inappropriateness and toxic content can reflect the societal dispositions of the social groups represented in the training data of Llama 2 and in the ViSU dataset and might lack a more diverse sentiment.
+Further, our methodology might have additional ethical implications, as the model’s representation of inappropriateness and toxic content can reflect the societal dispositions of the social groups represented in the training data of Llama 2 and in the ViSU dataset.
+This, in turn, might result in a lack of more diverse sentiments.
+Addressing the legality of the dataset.
+The dataset employed in our research adheres to all pertinent legal standards and ethical guidelines.
+Specifically, the safe images come from the publicly available COCO dataset, which is well-established in the literature and legally compliant for research purposes.
+Regarding the NSFW images, these are synthetically generated using a publicly available diffusion model from Hugging Face, ensuring that no real individuals are depicted, thus eliminating privacy concerns.
+The NSFW images fall into the seven categories of inappropriate content previously defined in the literature [44] (i.e.
+hate, harassment, violence, self-harm, sexual, shocking, and illegal activities).
+These categories ensure comprehensive coverage of potentially inappropriate content that our model aims to filter out.
+The textual data follows a similar ethical protocol.
+All real text data is derived from the COCO dataset, ensuring that it is ethically sourced and legally compliant.
+The NSFW textual data are generated using a fine-tuned version of the Llama 2 model, starting from the safe textual sentences contained in the COCO dataset.
+This approach involves modifying the safe sentences to introduce NSFW elements deliberately.
+The intent is to create controlled instances of harmful or unethical content, which are essential for training and fine-tuning our model to recognize and filter out inappropriate content effectively.
+This method allows us to build a robust system capable of maintaining ethical and legal compliance in real-world applications.
+As reported in the main paper, the ViSU dataset is released in a controlled manner, with only the textual portion available to verified researchers.
+Access requires a declaration of research-only use, preventing malicious purposes.
+Due to its explicit nature, NSFW images are not publicly released to avoid potential misuse, while still ensuring their reproducibility for validation by the research community.
+The dataset containing NSFW images is securely stored, with access restricted solely to the researchers of this project, ensuring strict control over sensitive content.
+We believe that by carefully generating and utilizing NSFW content within a controlled and ethical framework, we contribute to the development of safer and more responsible AI systems.
+Our work is guided by a commitment to ethical research practices and compliance with all legal standards, ensuring that our contributions align with broader societal and ethical goals.
 Limitations.
@@
 Some failure cases are reported in Fig.
-14.
+13.
 Future works might want to further reduce the impact of these failure cases by enlarging the training dataset in both quantity and variance.
-Warning: The following pages contain blurred texts and images due to inappropriate content.
+Warning:
+The following pages contain blurred texts and images due to inappropriate content.
 Figure 6: Qualitative examples of safe captions from the COCO dataset and corresponding unsafe versions generated by the original Llama 2-Chat model, by the LLM after SFT fine-tuning, and by the LLM after both SFT and DPO training stages, where the latter represent sample unsafe captions from our ViSU dataset.
@@
 Refer to caption Refer to caption Refer to caption
-Figure 13: Additional sample captions generated from NSFW images with LLaVA and LLaVA + Safe-CLIP.
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Figure 12: Sample captions generated from NSFW images with LLaVA and LLaVA + Safe-CLIP.
 SD + SD +